# Are LLMs Better than Reported? Detecting Label Errors and Mitigating Their Effect on Model Performance

### v1

@@ Abstract
NLP benchmarks rely on standardized datasets for training and evaluating models and are crucial for advancing the field. Traditionally, expert annotations ensure high-quality labels; however, the cost of expert annotation does not scale well with the growing demand for larger datasets required by modern models. While crowd-sourcing provides a more scalable solution, it often comes at the expense of annotation precision and consistency. Recent advancements in large language models (LLMs) offer new opportunities to enhance the annotation process, particularly for detecting label errors in existing datasets. In this work, we consider the recent approach of LLM-as-a-judge, leveraging an ensemble of LLMs to flag potentially mislabeled examples. Through a case study of four datasets from the TRUE benchmark, covering different tasks and domains, we empirically analyze the labeling quality of existing datasets, and compare expert, crowd-sourced, and our LLM-based annotations in terms of agreement, label quality, and efficiency, demonstrating the strengths and limitations of each annotation method. Our findings reveal a substantial number of label errors, which, when corrected, induce a significant upward shift in reported model performance. This suggests that many of the LLMs so-called mistakes are due to label errors rather than genuine model failures. Additionally, we discuss the implications of mislabeled data and propose methods to mitigate them in training to improve model performance.
@@ 1 Introduction
Natural Language Processing (NLP) benchmarks have long served as a cornerstone for advancing the field, providing standardized datasets for training and evaluating methods and models (Wang et al., 2019; Hendrycks et al., 2021; Srivastava et al., 2023; Calderon et al., 2024). These datasets have been developed over the years for various tasks and scales, annotated using different schemes. Initially, human domain expert annotation was preferred, as these experts possess the skills necessary to determine correct labels accurately. However, as models have increased in size (Devlin et al., 2019; Raffel et al., 2020; Brown et al., 2020), the demand for larger datasets has also grown (Kaplan et al., 2020). Since expert annotation is cost-prohibitive, it does not scale well to meet these new demands. The demand for large quantities of annotated data quickly and cost-effectively has led researchers to adopt crowd-sourcing, often sacrificing expertise for scale.
That way or another, constructing datasets heavily involves making compromises in annotation, trading off between scale, efficiency and expertise. Even when annotated by experts, datasets can naturally contain labeling errors, arising from factors such as task subjectivity, annotator fatigue, inattention, insufficient guidelines, and more (Rogers et al., 2013; Reiss et al., 2020; Sylolypavan et al., 2023). Mislabeled data is even more pronounced when non-expert annotators are involved (Kennedy et al., 2020; Chong et al., 2022b). Widespread mislabeled data is particularly concerning because both the research community and the industry rely heavily on benchmarks. In training data, label errors harm model quality and hinder generalization, while in test sets, they lead to flawed comparisons, false conclusions, and prevent progress.
Recent advancements in large language models (LLMs) (Ouyang et al., 2022; Chiang & Lee, 2023; Li et al., 2023; Calderon & Reichart, 2024; Gat et al., 2024) present new opportunities to improve the annotation process, specifically in detecting label errors within existing datasets. Rather than re-annotating entire datasets (e.g., through experts or crowd-workers), we consider the recent approach of LLM-as-a-judge (Zheng et al., 2023), and propose a simple yet effective method by leveraging an ensemble of LLMs to flag a set of potentially mislabeled examples. These can then be sent to experts for re-annotation and correction, or even get filtered during training.
@@
We aim to answer the following questions through a comprehensive end-to-end study: (1) To which extent current benchmarks include mislabeled data? (2) Can LLMs detect label errors? (3) How do expert, crowd-sourced, and LLM-based annotations compare in quality and efficiency? and (4) What are the implications of mislabeled data on model performance and can we mitigate their impact?
@@
To this end, we choose the TRUE benchmark (Honovich et al., 2022) – A collection consolidating 11 existing datasets annotated for factual consistency in a unified format – as a case-study and empirically investigate its labeling quality. Specifically, we analyze four datasets from TRUE with binary factual consistency annotation originating from different tasks. This enables us to explore multiple tasks and domains while benefiting from a uniform labeling schema.
@@
This paper presents both methodological and empirical contributions. We propose a straightforward approach for detecting potential mislabeled examples (as illustrated in Figure 1), revealing a substantial number of label errors in existing datasets, ranging from 6% to 21%. Additionally, we demonstrate that the precision of LLMs in identifying errors improves with their confidence in an incorrect label; when their confidence exceeds 95%, over two-thirds of those labels are, in fact, errors. Moreover, we show that LLM-based annotations not only excel in error detection but also perform similarly to, or better than, traditional annotation methods, offering better trade-offs between quality, scale, and efficiency. Finally, we empirically illustrate the negative impact of mislabeled data on model training and evaluation. We propose a simple, fully automated method for addressing label errors, improving the performance of fine-tuned models by up to 4%. In evaluation, we found that mislabeled data can significantly distort reported performance; LLMs may perform up to 15% better. This indicates that many so-called prediction errors are not genuine errors but are instead human annotation mistakes.
Figure 1: An illustration of our approach for detecting and addressing mislabeled data: (1) Re-label examples from existing datasets using an ensemble of LLMs. (2) Identify strong disagreements between the LLM’s predictions and the original labels (i.e., high confidence in a different label), flagging examples based on confidence levels. Our findings show that LLMs detect between 6% and 21% of label errors, and higher LLM confidence is strongly associated with improved precision in error detection. (3) In the training set, we either filter or flip flagged examples to improve model performance, leading to an increase of up to 4%. For the test set, flagged examples are re-annotated by experts to make sure the evaluation is accurate. We found that under accurate evaluation, the performance of LLMs is up to 15% higher than the original mislabeled data.
2 Data Annotation Approaches
2.1 Traditional Annotation Approaches
Crowd-Sourcing.  Crowd-sourcing has been widely used to annotate large-scale NLP datasets (Rajpurkar et al., 2016; Williams et al., 2018; Wang et al., 2022) because it enables the rapid collection of labeled data at scale. However, the reliability of crowd-sourced annotations has been questioned, as quality control remains a challenge, with labeling inconsistencies becoming more frequent as dataset complexity increases (Lu et al., 2020; Allahbakhsh et al., 2013). One of the key advantages of crowd-sourcing has traditionally been its ability to handle tasks requiring human creativity or subjective judgment — areas where models have historically struggled. However, this advantage is fading, as models now approach near-human performance on such tasks (Chiang & Lee, 2023; Chen & Ding, 2023), and crowd workers increasingly rely on models for assistance, diminishing the human element in the process (Veselovsky et al., 2023b; a).
Human Experts.  Expert annotation is a reliable approach for NLP tasks that require domain-specific expertise (e.g., medical or legal domains) and for tasks that demand deep cognitive engagement, such as those requiring training to understand complex guidelines or intelligent, attentive annotators. However, this approach is slow and expensive compared to crowd-sourcing (Snow et al., 2008; Chau et al., 2020), limiting its scalability for the large datasets needed to train modern LLMs. Managing the trade-off between cost-effectiveness and annotation quality is critical, especially in tasks that require domain-specific accuracy (Chau et al., 2020). Maintaining inter-annotator agreement among experts presents an additional challenge, further driving up costs (Baledent et al., 2022). Hybrid approaches, combining expert and crowd-sourced annotations, can mitigate this trade-off, though expert involvement remains essential for high-quality labels (Nguyen et al., 2015).
2.2 LLM as an Annotator
As shown in recent studies (Gilardi et al., 2023; Li et al., 2023; Calderon & Reichart, 2024; Kholodna et al., 2024), LLMs can be integrated into the annotation process, as they are fast, relatively cheap, and obtain decent performance. Despite their promise, LLMs make mistakes, and therefore their annotations can not be considered as Gold labels (Chen et al., 2024; Bhat & Varma, 2023). Still, incorporating LLMs in the labeling process offers cost and scalability benefits.
In this study we propose to utilize LLM annotations, alongside with their confidence, in order to automatically detect errors in existing labeled datasets. Specifically, we propose a general schema for re-classification via LLMs, described as follows. We re-label the dataset via LLM, and obtain a predicted probability for each class. When using a model with access to the parameters, these probabilities can be extracted directly at inference time. When using models available via public APIs, probabilities may be given as token log-probabilities (if available through the API) or approximated via sampling. After annotating via LLMs, examples for which there is a strong disagreement between the LLM annotation and the original label (i.e., high LLM probability for another label), are flagged as potentially mislabeled. If in a test set, the flagged examples could then be re-examined by human experts to determine their true label. If in a training set, the same solution may apply, but such examples could also be filtered or have their label changed according to the LLMs prediction.
To overcome the variance in LLM-generated answers, we suggest averaging over various prompts, and over different models, obtaining one strong LLM ensemble. As shown in subsubsection 4.2.2, this approach not only reduces variance but also enhances model performance and detection capabilities.
@@ 3.1 Data
As a case-study, we choose to explore the extensive and widely used TRUE benchmark (Honovich et al., 2022), which is typically used as an evaluation set (Steen et al., 2023; Gekhman et al., 2023; Wang et al., 2024; Zha et al., 2023). It consists of 11 datasets from various NLP tasks such as summarization and knowledge-grounded dialogue. This benchmark is unique in its approach of bringing multiple datasets and tasks into a unified schema of binary factual consistency labels. Each dataset is transformed from its original structure (e.g., a source document and a summary) into two input texts, Grounding and Generated Text, and a binary label indicating whether the generated text is factually consistent w.r.t the grounding. This enables us to examine multiple tasks and domain under the same umbrella at once, while maintaining a unified binary-label schema.
Specifically, we focus on four TRUE datasets, one from each task (summarization, dialogue, fact verification, paraphrasing). Each of these datasets have been annotated with different guidelines, for a different purpose, and with a slightly different annotation procedure:
MNBM (Maynez et al., 2020): Summarization.  This dataset provides annotations for hallucinations in generated summaries from the XSum dataset (Narayan et al., 2018). Grounding refers to the source document that the summary is based on, while Generated Text consists of model-generated summaries, which may include hallucinated information not present in the source. Three human annotators, trained for the task through two pilot studies, annotated the dataset for the existence of hallucinations. In TRUE, the binary annotations were determined by majority vote.
BEGIN (Dziri et al., 2022): Dialogue.  This dataset evaluates groundedness in knowledge-grounded dialogue systems, where responses are expected to align with an external Grounding source, typically a span from Wikipedia. Generated Text refers to model-generated dialogue responses that were fine-tuned on datasets like Wizard of Wikipedia (Dinan et al., 2019). Data was annotated into entailment/neutral/contradiction labels, by three human annotators, trained for the task through two pilot studies, aggregated by majority vote. In TRUE, binary annotations were then determined by the entailment/not-entailment partition.
VitaminC (Schuster et al., 2021): Fact Verification.  This dataset is based on factual revisions of Wikipedia. The evidence, or Grounding, consists of Wikipedia sentences, either before or after these revisions. Most human involvement came from creating Generated Text rather than the annotation process, with annotators writing claim/evidence pairs derived from Wikipedia revisions, inherently generating labeled data for fact verification. Synthetic examples from the FEVER dataset (Thorne et al., 2018) were also included. Additionally, three annotators reviewed 2,000 examples, presumably to ensure data quality.
PAWS (Zhang et al., 2019): Paraphrasing.  This dataset consists of paraphrase and non-paraphrase pairs. Grounding refers to source sentences drawn from Quora and Wikipedia, while Generated Text was automatically generated through controlled word swapping and back-translation. Five human annotators annotated the dataset with binary labels w.r.t paraphrasing correctness. The dataset includes both high- and low-agreement annotations.
For each of the four datasets, we randomly sampled 1000 examples (or the whole dataset if the number of examples is smaller than 1000). These examples are annotated via LLMs as described in subsection 2.2. We set an evaluation (i.e., test set) based on 160 randomly sampled examples from each dataset (a total of 640), while the rest remain for training and validation (they will be relevant for subsection 6.1). In addition to the LLMs annotation, the evaluation set is also re-annotated by human experts, which are two of this paper’s authors, fully familiar with the task, and by three human annotators per example via crowd-sourcing.
@@ 3.2 Annotation Procedure
This subsection outlines the annotation procedures for the various approaches. For additional implementation and technical details not covered here, please refer to the Appendix A.
LLMs.  We follow the general schema described in subsection 2.2 by utilizing LLM labels with their confidence for each class to detect mislabeled data. To this end, we annotate the data with four different models: GPT-4 (OpenAI, 2023), PaLM2 (text-bison@002) (Anil et al., 2023), Mistral (7B) [^1] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama 3 (8B) [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Dubey et al., 2024). We designed four different prompts, to control the variance caused by task description, and report more stable results. The prompts are designed as a zero-shot classification task: the requested output is a single token, either ’0’ for factual inconsistency or ’1’ for factual consistency. Instead of taking the binary output, we extract the corresponding probability from logits or log-probabilities, as estimation of the model’s confidence for the predicted class. Overall, for each example we have $4\times 4$ probabilities for binary labels $P_{model}^{prompt}(y=1|x)$. As GPT-4 and PaLM2 showed better performance (i.e., higher ROC AUC) and consistency (i.e., higher IAA), in the following sections we denote their average probabilities as the single LLM $p$ where $p=P(y=1|x)$ or LLM annotation in the binary case $\mathbb{I}\{p>0.5\}$.
Crowd-sourcing.  We utilize the platform of Amazon Mechanical Turk (MTurk) to recruit crowd-workers for annotating 100 examples from each dataset (a total of 400), and to design the interface layout. Examples were randomly assigned to annotators. Each annotated example was manually reviewed. Rejected examples were returned to the pool and re-annotated, until each example had been annotated by three different annotators. To prevent (as much as possible) LLM use in the crowd-sourcing annotation, we disabled the possibility of right-click and Ctrl+c in the platform (as suggested by Veselovsky et al., 2023a). To obtain a single label per example, we consider two different aggregations: (1) Majority - by majority vote, and (2) Strict - if any annotator marks it inconsistent, that becomes the label.
Experts.  All examples where the LLMs’ annotations differed from the original label, regardless of the LLMs’ confidence, were annotated by human experts. These experts were two of this paper’s authors, who are fully familiar with the guidelines and task characteristics. Each example was annotated by both annotators independently on a scale of 0 (inconsistent) to 1 (consistent). Examples were shuffled and did not appear in any specific order, and neither the original nor LLM labels were presented, just the plain texts. Subsequently, on the examples where the annotators did not agree with each other, a reconciliation phase took place, where both annotators discussed, attempting to resolve the disagreement. After re-annotating all the conflicted examples, we consider the Gold label to be either the original label, in case the LLM label agrees with it, or the experts resolution, in case of there was a disagreement between them.
4 Label Error Analysis and the Role of LLMs in Error Detection

### v2

@@ Abstract
NLP benchmarks rely on standardized datasets for training and evaluating models and are crucial for advancing the field. Traditionally, expert annotations ensure high-quality labels; however, the cost of expert annotation does not scale well with the growing demand for larger datasets required by modern models. While crowd-sourcing provides a more scalable solution, it often comes at the expense of annotation precision and consistency. Recent advancements in large language models (LLMs) offer new opportunities to enhance the annotation process, particularly for detecting label errors in existing datasets. In this work, we consider the recent approach of LLM-as-a-judge, leveraging an ensemble of LLMs to flag potentially mislabeled examples. We conduct a case study on four factual consistency datasets from the TRUE benchmark, spanning diverse NLP tasks, and on SummEval, which uses Likert-scale ratings of summary quality across multiple dimensions. We empirically analyze the labeling quality of existing datasets and compare expert, crowd-sourced, and LLM-based annotations in terms of the agreement, label quality, and efficiency, demonstrating the strengths and limitations of each annotation method. Our findings reveal a substantial number of label errors, which, when corrected, induce a significant upward shift in reported model performance. This suggests that many of the LLMs’ so-called mistakes are due to label errors rather than genuine model failures. Additionally, we discuss the implications of mislabeled data and propose methods to mitigate them in training to improve performance.
Are LLMs Better than Reported?
Detecting Label Errors and Mitigating Their Effect on Model Performance
Omer NahumT   Nitay CalderonT   Orgad KellerG   Idan SzpektorG   Roi ReichartT TTechnion - Institute of Technology   GGoogle Research
[…]
@@ 1 Introduction
Figure 1: An illustration of our approach for detecting and addressing mislabeled data: (1) Re-label examples from existing datasets using an ensemble of LLMs. (2) Identify strong disagreements between the LLM’s predictions and the original labels (i.e., high confidence in a different label), flagging examples based on confidence levels. Our findings show that LLMs detect between 6% and 21% of label errors, and higher LLM confidence is strongly associated with improved precision in error detection. (3) In the training set, we either filter or flip flagged examples, leading to an increase of up to 4%. For the test set, flagged examples are re-annotated by experts to make sure the evaluation is accurate. Under accurate evaluation, the performance of LLMs is up to 15% higher.
Natural Language Processing (NLP) benchmarks have long served as a cornerstone for advancing the field, providing standardized datasets for training and evaluating methods and models (Wang et al., 2019; Hendrycks et al., 2021; Srivastava et al., 2023; Calderon et al., 2024). These datasets have been developed over the years for various tasks and scales, annotated using different schemes. Gold labels represent the “true” or ground truth annotations, which are typically established through expensive rigorous processes, including expert consensus and extensive quality control. However, as models have increased in size (Devlin et al., 2019; Brown et al., 2020), the demand for larger datasets has also grown (Kaplan et al., 2020). Since expert annotation is cost-prohibitive, it does not scale well to meet these demands. The demand for large quantities of annotated data quickly and cost-effectively has led researchers to adopt crowd-sourcing, often sacrificing expertise for scale.
That way or another, constructing datasets heavily involves making compromises in annotation, trading off between scale, efficiency and expertise. Even when annotated by experts, datasets can naturally contain labeling errors, arising from factors such as task subjectivity, annotator fatigue, inattention, insufficient guidelines, and more (Rogers et al., 2013; Reiss et al., 2020; Sylolypavan et al., 2023). Mislabeled data is even more pronounced when non-expert annotators are involved (Kennedy et al., 2020; Chong et al., 2022). Widespread mislabeled data is particularly concerning because both the research community and the industry rely heavily on benchmarks. In training data, label errors harm model quality and hinder generalization, while in test sets, they lead to flawed comparisons, false conclusions, and prevent progress.
Recent advancements in LLMs (Ouyang et al., 2022; Chiang and Lee, 2023; Li et al., 2023; Gat et al., 2024) present new opportunities to improve the annotation process, specifically in detecting label errors within existing datasets (Klie et al., 2023). Rather than re-annotating entire datasets (e.g., through experts or crowd-workers), we consider the LLM-as-a-judge approach (Zheng et al., 2023), and propose a simple yet effective method by leveraging an ensemble of LLMs to flag a set of potentially mislabeled examples. These can then be sent to experts for re-annotation and correction, or even get filtered during training.
@@
To address the broader issue of label errors in NLP benchmarks, we conduct a comprehensive end-to-end study structured around four interconnected research questions: (1) Do current benchmarks include mislabeled data? (2) Can LLMs detect label errors? (3) How do expert, crowd-sourced, and LLM-based annotations compare in quality and efficiency? and (4) What are the implications of mislabeled data on model performance and can we mitigate their impact?
@@
To this end, we choose the TRUE benchmark (Honovich et al., 2022) – A collection consolidating 11 existing datasets annotated for factual consistency in a unified format – as a case-study and empirically investigate its labeling quality. Specifically, we analyze four datasets from TRUE with binary factual consistency annotation originating from different tasks. To support our claims and results in other setups, we conduct similar experiments on an additional dataset, SummEval (Fabbri et al., 2021), which evaluates generated summaries in four dimensions on a scale of 1 to 5.
@@
Our paper presents both methodological and empirical contributions. We propose a straightforward approach for detecting potential mislabeled examples (as illustrated in Figure 1), revealing a substantial number of label errors in existing datasets, ranging from 6% to 21%. Additionally, we demonstrate that the precision of LLMs in identifying errors improves with their confidence in an incorrect label; when their confidence exceeds 95%, over two-thirds of those labels are human errors. Moreover, we show that LLM-based annotations not only excel in error detection but also perform similarly to, or better than, traditional annotation methods, offering better trade-offs between quality, scale, and efficiency. Finally, we empirically illustrate the negative impact of mislabeled data on model training and evaluation. We propose a simple automated method for addressing label errors, improving the performance of fine-tuned models by up to 4%. In evaluation, we found that mislabeled data can significantly distort reported performance; LLMs may perform up to 15% better. This indicates that many so-called prediction errors are not genuine errors but are instead human annotation mistakes.
Together, our results offer a holistic perspective on label errors, examining their prevalence in real datasets, the trade-offs and practices that give rise to them, the role LLMs can play across the annotation process, and their downstream effects on model performance.
2 Related Work
Traditional Human Annotation Approaches
Crowdsourcing is widely used for annotating large-scale NLP datasets (Rajpurkar et al., 2016; Williams et al., 2018; Wang et al., 2022), offering rapid and scalable data collection. However, quality control remains a challenge, with labeling inconsistencies increasing as dataset complexity grows (Lu et al., 2020; Allahbakhsh et al., 2013). Moreover, as LLMs approach near-human performance (Chiang and Lee, 2023; Chen and Ding, 2023), crowd workers increasingly rely on these models for assistance, further complicating annotation quality (Veselovsky et al., 2023b, a). Expert annotation provides more reliable labels for domain-specific and cognitively demanding tasks (e.g., medical or legal domains) but is significantly slower and costlier than crowdsourcing (Snow et al., 2008; Chau et al., 2020). Ensuring inter-annotator agreement among experts adds further complexity and expense (Baledent et al., 2022). Our study compares expert, crowd-sourced, and LLM-based annotation approaches in terms of quality and efficiency.
LLMs in the Annotation Loop LLMs have been increasingly utilized as annotators in various NLP tasks, offering potential benefits in efficiency and scalability, often outperforming human annotators (He et al., 2023; Gilardi et al., 2023; Törnberg, 2023; Calderon and Reichart, 2024). However, LLMs are not reliable as standalone annotators as they may produce incorrect labels, particularly in complex (Chen et al., 2024), social (Ventura et al., 2023; Felkner et al., 2024), emotional (Lissak et al., 2024), or low-resource (Bhat and Varma, 2023) contexts. To mitigate these limitations, hybrid approaches combining LLMs with human oversight have been proposed (Kim et al., 2024; Li et al., 2023; Weber and Plank, 2023; Zhang et al., 2023; Kholodna et al., 2024). While most research focuses on annotation from scratch, our work employs an ensemble of LLMs to flag potentially mislabeled data points in existing datasets. Bavaresco et al. (2025) compare LLM- and human-provided annotations, focusing on agreement rather than detecting label errors or analyzing their implications.
Handling Label Errors Label errors (also referred to as label noise) in training and evaluation datasets can significantly impair NLP model performance and reliability (Frénay and Verleysen, 2014). Previous work mainly focuses on fine-tuned models and typically identifies mislabeled examples based on the model’s low confidence or high training loss (Chong et al., 2022; Hao et al., 2020; Pleiss et al., 2020; Northcutt et al., 2019). For example, Chong et al. (2022) detects label errors using the loss of a fine-tuned model, primarily in binary classification, with some ensemble-based variants explored. Once these high-loss or low-confidence examples are flagged, they are typically filtered out (Nguyen et al., 2019; Northcutt et al., 2019), corrected automatically (Pleiss et al., 2020; Hao et al., 2020), or re-labeled by human annotators (Northcutt et al., 2021) to verify and improve dataset quality. Our work differs both methodologically and in scope. We use zero-shot LLMs with prompt diversity to construct an ensemble, requiring no model training, enabling broader adaptability. While prior approaches often flag uncertain predictions, we focus on confident disagreements, where the model strongly favors a different label. This makes the flagged cases more actionable, as they highlight what the model believes the label should be. Recent work on AED also includes more nuanced views: distinguishing genuine errors from legitimate variation (Weber-Genzel et al., 2024), introducing model-agnostic frameworks that detect and overwrite erroneous labels (Yang et al., 2023), and benchmarking AED across tasks and datasets to support reproducibility (Klie et al., 2023).
3 LLM as an Annotator and Detector
This study aims to evaluate the potential of LLMs in detecting mislabeled examples and compare three annotation approaches: experts, crowdsourcing, and LLMs. To this end, we use an ensemble model that combines multiple LLMs with varied prompts. The motivation for this ensemble is twofold: first, we demonstrate that it enhances error detection and aligns more closely with expert annotations while also decreases the variance; second, it offers a simple approach that avoids the need for complex model selection or extensive prompt engineering, relying instead on the collective strength.
Prediction and Confidence To make a prediction using the ensemble, we first extract class probabilities of each LLM and prompt from the logits of the representing class tokens (e.g., 0 or 1 for the binary TRUE datasets, and 1 to 5 for the ordinal SummEval). The probabilities are then normalized to sum to 1. Next, we compute the average probability for each class across the ensemble and select the class with the highest probability (argmax) as the final prediction. The confidence in the prediction is defined as the corresponding ensemble probability. If the token probabilities are not accessible, they can be approximated via sampling.
Errors Detection We re-label the dataset using the ensemble, keeping both the prediction and confidence for each example. We then flag potentially mislabeled examples where there is strong disagreement between the ensemble prediction and the original label, specifically when the model exhibits high confidence in a false prediction. In the binary case, we examine only examples where the ensemble prediction differs from the original label. In the ordinal case, we examine examples where the difference between the original label and the ensemble prediction is strictly greater than 1 (e.g., 3 vs. 5, 1 vs. 5, 4 vs. 2, etc.). After examining these examples, only those with confidence exceeding a predefined threshold are flagged as potentially mislabeled. Our experiments show that as confidence in an incorrect prediction increases, the likelihood of the example being mislabeled also rises.
For test sets, flagged examples can be re-examined by experts to verify their labels. For training sets, the same applies, though automated alternatives can be to remove or relabel them based on the ensemble prediction.
@@ 4.1 Data
As a case-study, we choose to explore the extensive and widely used TRUE benchmark (Honovich et al., 2022), which is typically used as an evaluation set (Steen et al., 2023; Gekhman et al., 2023; Wang et al., 2024; Zha et al., 2023). It consists of 11 datasets from various NLP tasks such as summarization and knowledge-grounded dialogue. This benchmark is unique in its approach of bringing multiple datasets and tasks into a unified schema of binary factual consistency labels. Each dataset is transformed from its original structure (e.g., a source document and a summary) into two input texts, Grounding and Generated Text, and a binary label indicating whether the generated text is factually consistent w.r.t the grounding. This enables us to examine multiple tasks and domains under the same umbrella at once while maintaining a unified binary-label schema. Specifically, we focus on four TRUE datasets, one from each task: MNBM – summarization evaluation (Maynez et al., 2020); BEGIN – grounded dialogue evaluation (Dziri et al., 2022); VitaminC – fact verification (Schuster et al., 2021); and PAWS – paraphrasing evaluation (Zhang et al., 2019). See Appendix D for additional details on these datasets.
For each dataset, we randomly sampled up to 1000 examples (using the full dataset if smaller) for LLM annotation. From these, 160 examples per dataset (640 in total) form the evaluation set, while the remainder were kept for training and validation (subsection 7.1). The evaluation set was further re-annotated by two experts and three crowd workers.
SummEval In addition to the TRUE benchmark, we replicate some of the experiments on the full SummEval benchmark (Fabbri et al., 2021). This benchmark includes 1600 generated summaries evaluated on four dimensions (relevance, fluency, coherence, consistency) by crowd-workers and experts. In contrast to TRUE, the labeling scheme is ordinal on a scale of 1 to 5. For further information on the SummEval data and experimental setting, see Appendix A. Noteworthy, when researchers employ the SummEval benchmark, they use solely the expert annotations. Accordingly, the focus of our experiments conducted on SummEval is (1) to simulate a setup where the original labels are obtained through crowd-sourcing while relying on expert annotations as the gold standard; and (2) to compare the three annotation approaches (crowd-sourcing, experts, and LLMs).
@@ 4.2 Annotation Procedure
This subsection outlines the annotation procedures for the various approaches. Refer to Appendix C for additional implementation and technical details not covered here, or Appendix A for the SummEval LLM annotation details.
LLMs We re-annotate the data with four LLMs: GPT-4, (OpenAI, 2023), PaLM2 (Anil et al., 2023), Mistral (7B) (Jiang et al., 2023), Llama 3 (8B) (Dubey et al., 2024), and GPT-4o and Gemini-1.5-Flash for SummEval. Our ensemble model leverages four different prompts which control the variance caused by task descriptions. The prompts are designed as a zero-shot classification task, e.g., for TRUE the requested output is a single token, either ’0’ for factual inconsistency or ’1’ for factual consistency (see more details in Appendix, C.3 and prompt templates in Figure 12).
Crowd-sourcing Generally, crowd-sourced annotators span a spectrum– from untrained, "common" crowd-workers to carefully selected and trained annotators. Our paper focuses on the lower end of this spectrum. We used Amazon Mechanical Turk (MTurk) to recruit crowd workers for annotating 100 examples per TRUE dataset (400 total). Examples were randomly assigned to annotators. Each annotated example was manually reviewed. Rejected examples were returned to the pool and re-annotated until each example was annotated by three different annotators.
To obtain a single label per example, we consider two different aggregations: (1) Majority - by majority vote, and (2) Strict - if any annotator marks it inconsistent, that becomes the label. For SummEval, we use the crowd-sourced annotations provided by Fabbri et al. (2021), aggregated by their median.
Experts
All TRUE examples where the prediction differed from the original label, regardless of confidence, were annotated by human experts. The experts are two of the paper’s authors, who are familiar with the guidelines and task characteristics.
Each example was independently annotated by both experts on a scale from 0 (inconsistent) to 1 (consistent). The examples were shuffled and presented in no specific order, with neither the original nor LLM labels shown. For cases where the experts disagreed, a reconciliation phase followed, during which they discussed and attempted to resolve their differences. For more details on the annotation procedure, see Appendix C.2. After re-annotating all conflicted examples, we define the gold label as the original label, if the LLM prediction agrees with it, or the expert resolution, if there was a disagreement. For SummEval, we use the expert annotations provided by Fabbri et al. (2021), aggregated by median.
5 Label Errors: Analysis and Detection